JET-COOLED Ã-X̃ SPECTRA OF THE β-HYDROXYETHYLPEROXY and β-HYDROXYETHYLPEROXY-OD RADICALS

\underline{MING-WEI CHEN}, GABRIEL M. P. JUST^{a}, TERRANCE CODD, and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210.

The β-hydroxyethylperoxy (HOCH_{2}CH_{2}OO, β-HEP) radical in the atmosphere arises from the reaction of ethene (CH_{2}CH_{2}) and hydroxyl radical (OH), followed by the reaction with oxygen (O_{2}). It is also an important intermediate in the oxidation of ethanol, a component of automotive fuel. High-resolution, jet-cooled cavity ring-down spectroscopy (resolution of Δν ≈250MHz, considering the instrumental linewidth and the residual Doppler broadening) has been applied to observe the Ã-X̃ origin band of the most stable conformer of both β-HEP and mono-deuterated β-HEP (DOCH_{2}CH_{2}OO, β-HEP-OD). Broadened rotational contours are observed for both isotopologues, but more resolved structure is shown in the β-HEP-OD spectrum. The evolutionary algorithm approach is applied to analyze the spectra, which gives fitted rotational constants and the homogeneous linewidths for both isotopologues using an asymmetric-top model for the rotational Hamiltonian. Evidence corresponding to a narrower homogeneous linewidth in the β-HEP-OD spectra implies that the broad rotational contour of β-HEP likely involves the motion of the hydrogen of the OH group.

apresent address: Lawrence Berkeley National Laboratory, Lawrence Berkeley National Laboratory